# Rotatable Antenna Enabled Wireless Communication: Modeling and Optimization

### v4

@@ II-B Channel Model
where $\mathbf{F}\triangleq[\vec{\mathbf{f}}_{1},\vec{\mathbf{f}}_{2},\dots,\vec{\mathbf{f}}_{N}]\in\mathbb{R}^{3\times N}$ is the pointing matrix of all RAs, $\mathbf{h}_{k}^{\mathrm{LoS}}(\mathbf{F})\triangleq[h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{1}),h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{2}),\dots,h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{N})]^{T}$ and $\mathbf{h}_{k}^{\mathrm{NLoS}}(\mathbf{F})\triangleq[h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{1}),h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{2}),\dots,h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{N})]^{T}$ are the LoS and NLoS channel components from the BS to user $k$, respectively. It is observed that variations in $\mathbf{F}$ lead to the rotation of the directional gain pattern of each RA, thereby altering the effective multipath channel in (10). Note that the current geometric channel model can also be extended to incorporate more realistic propagation characteristics by adopting the standardized path loss models defined in 3GPP TR 38.901 [33].
@@
Remark 1: Although antenna orientation/boresight adjustments inevitably alter antenna polarization characteristics in practice, the proposed RA architecture introduces only minor variations in polarization characteristics due to its constrained rotational range at fixed antenna position and the absence of self-rotation. Moreover, the polarization ellipticity and handedness remain constant throughout the adjustment process, and thus efficient polarforming techniques [34, 35] can be applied to compensate the polarization induced channel variation such that the orientation/boresight adjustment by RA is still effective. Therefore, to focus on the effects of the directional gain pattern and explore the spatial DoFs offered by antenna orientation/boresight adjustments, polarization modeling is not considered in this paper for simplicity. Nevertheless, polarization modeling can be incorporated into future work on RA by integrating polarization matching efficiency into the channel power gain, following approaches similar to [34] and [35]. Such extensions would require joint consideration of both the directional gain pattern and polarization matching during orientation/boresight optimization.
@@ II-C Min-SINR Maximization Problem
In this paper, we aim to maximize the minimum SINR among all the users by jointly optimizing the receive beamforming matrix $\mathbf{V}\triangleq[\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{K}]$ and RA’s pointing matrix $\mathbf{F}$, subject to the rotational constraint given in (4). Thus, the min-SINR maximization problem is formulated as
[…]
@@
where constraint (14b) is equivalent to (4) for ensuring that the rotation of RAs’ boresight directions does not exceed the given range, and constraint (14c) ensures that $\vec{\mathbf{f}}_{n}$ is a unit vector.
In practice, channel estimation is necessary at the BS to obtain the required channel state information (CSI). Since adjusting the antenna’s orientation/boresight rather than its position does not change the propagation geometry, as shown in (8) and (9), conventional MIMO channel estimation methods for fixed antennas remain applicable to the RA system [22]. Furthermore, the flexible adjustments of RAs’ orientations/boresights enable the BS to capture pilot signals from diverse directions of arrival, yielding richer channel information and enhanced CSI estimation accuracy.
Denoting $\hat{\mathbf{h}}_{k}(\mathbf{F})$ as the estimated CSI of the link from user $k$ to the BS, the received signal is expressed as
@@
where $\check{\mathbf{h}}_{k}\triangleq\mathbf{h}_{k}(\mathbf{F})-\hat{\mathbf{h}}_{k}(\mathbf{F})$ stands for the CSI estimation error that is uncorrelated to $\mathbf{h}_{k}(\mathbf{F})$. If the CSI estimation error follows the zero-mean complex Gaussian distribution, i.e., $\sum_{k=1}^{K}{\check{\mathbf{h}}_{k}\sqrt{P_{k}}s_{k}}\sim\mathcal{N}_{c}\left(\mathbf{0},e^{2}\mathbf{I}_{N}\right)$ with $e$ accounting for the level of channel estimation error [36], the resultant error term can be incorporated into the AWGN term in (15) for simplicity. Accordingly, the receive SINR for user $k$ is given by
@@ III Single-User Case with Free-Space Propagation
[…]
@@ III-A Optimal Closed-Form Solution
For any given RA’s pointing matrix $\mathbf{F}$ in the single-user case, it is known that the MRC beamformer is the optimal receive beamforming solution to problem (P2) [37], i.e., $\mathbf{v}_{\mathrm{MRC}}=\frac{\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})}{\|\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})\|}$. Thus, substituting $\mathbf{v}_{\mathrm{MRC}}$ into (17a) yields the following SNR expression,
$\gamma=\bar{P}\|\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})\|^{2}=\frac{\bar{P}A}{4\pi}\sum_{n=1}^{N}{\frac{G_{0}\operatorname{cos}^{2p}(\epsilon_{n})}{r_{n}^{2}}}.$ (18)
Exploiting the structure in (18), problem (P2) can be decomposed into $N$ subproblems, each of which independently optimizes the pointing vector of one RA. For RA $n$, the corresponding subproblem is given by
[…]
@@
[…]
@@ III-B Asymptotic Performance Analysis
In this subsection, we focus on the performance analysis for the single-user system, and derive a closed-form expression for the SNR in the ULA case, lower/upper bounds for the SNR in the UPA case, and the asymptotic gains in both cases as the antenna number $N$ goes to infinity. For ease of exposition, we assume that the user is located along the $z$-axis (i.e., $\psi=\frac{\pi}{2}$ and $\phi=0$) and its position is denoted by $\mathbf{u}=[0,0,r]^{T}$. In this case, based on the optimal pointing vector obtained in (20), the entire array region can be divided into inner and outer areas by comparing $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})$ and $\theta_{\mathrm{max}}$, as shown in Fig. 3. Specifically, if RA $n$ is located in the inner area, we have $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})\leq\theta_{\mathrm{max}}$ and $\epsilon_{n}=0$, i.e., RA $n$ can adjust its boresight to perfectly align with the user direction to obtain the maximum antenna directional gain. Conversely, for RA $n$ located in the outer area, we have $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})>\theta_{\mathrm{max}}$ and $\epsilon_{n}=\operatorname{arctan}\left(\sqrt{(n_{x}^{2}+n_{y}^{2})\delta^{2}}\right)-\theta_{\mathrm{max}}$, i.e., RA $n$ can only serve the user with $\theta_{\mathrm{z},n}=\theta_{\mathrm{max}}$ and the RA boresight is offset from the user direction by an angle $\epsilon_{n}>0$ due to the zenith angle constraint in (4). Based on the above discussion, the projection between the user direction vector and the optimal pointing vector of RA $n$ can be expressed as
$\operatorname{cos}\epsilon_{n}=\operatorname{cos}\left(\left[\operatorname{arctan}\left(\sqrt{(n_{x}^{2}+n_{y}^{2})\delta^{2}}\right)-\theta_{\mathrm{max}}\right]^{+}\right),$ (22)
where $[x]^{+}\triangleq\mathrm{max}\{0,x\}$. By substituting (22) into (18), we can obtain the resultant SNR as (23), shown at the top of next page. The SNR in (23) involves a double summation, which may make it difficult to gain useful insights. By approximating the double summation in (23) as its corresponding double integral by leveraging $\delta\ll 1$ as in [38, 39, 40, 41], the SNR can be rewritten in an integral form as (24), also shown at the top of next page.
@@
$\gamma=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi}\sum_{n_{x}=-\frac{N_{x}-1}{2}}^{\frac{N_{x}-1}{2}}{\sum_{n_{y}=-\frac{N_{y}-1}{2}}^{\frac{N_{y}-1}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(\sqrt{(n_{x}^{2}+n_{y}^{2})\delta^{2}}\right)-\theta_{\mathrm{max}}\right]^{+}\right)}{1+(n_{x}^{2}+n_{y}^{2})\delta^{2}}}}.$ (23)
$\gamma\simeq\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi\Delta^{2}}\int_{-\frac{N_{y}\Delta}{2}}^{\frac{N_{y}\Delta}{2}}{\int_{-\frac{N_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(\sqrt{\frac{1}{r^{2}}(x^{2}+y^{2})}\right)-\theta_{\mathrm{max}}\right]^{+}\right)}{1+\frac{1}{r^{2}}(x^{2}+y^{2})}\mathrm{d}x\mathrm{d}y}}.$ (24)
1) ULA-Based RA System: To gain some insights, we first focus on the ULA setting. With $N_{y}=1$ and $N=N_{x}$, the resultant SNR in (24) reduces to
$\gamma=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi\Delta}\int_{-\frac{N_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(|\frac{x}{r}|\right)-\theta_{\mathrm{max}}\right]^{+}\right)}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}.$ (25)
@@
${\gamma}=\begin{cases}{\frac{2\bar{P}\xi\delta}{\pi}\triangle_{\mathrm{span}}(N_{x}),}&{N_{x}\leq\bar{N}_{x}}\\ {\frac{2\bar{P}\xi\delta}{\pi}\left[\theta_{\mathrm{max}}+\operatorname{sin}\left(\triangle_{\mathrm{span}}(N_{x})-\theta_{\mathrm{max}}\right)\right],}&{N_{x}>\bar{N}_{x},}\end{cases}$ (26)
@@
Theorem 1 shows that with the applied MRC receive beamforming and the optimized RAs’ pointing vectors, the resultant maximum SNR of the proposed ULA-based RA system scales with the antenna number $N_{x}$ according to the span angle function $\triangle_{\mathrm{span}}(N_{x})$. Furthermore, given the antenna size $A$, the antenna separation $\Delta$, and the propagation distance $r$, the maximum SNR of the ULA-based RA system mainly depends on the allowable range for antenna boresight rotation and the ULA size.
Remark 2: By applying the linear approximation for the arctangent function, i.e., $\operatorname{arctan}(x)\approx\frac{\pi}{4}x,\;-1\leq x\leq 1$ [42], the SNR obtained in the first case of (26) can be approximated by $\gamma\approx\frac{1}{2}N_{x}\bar{P}\xi\delta$ since we have $0\leq\frac{N_{x}\delta}{2}\leq 1$ when $N_{x}\leq\bar{N}_{x}$. Thus, the resultant SNR increases linearly with the number of RAs when $N_{x}\leq\bar{N}_{x}$, i.e., $\triangle_{\mathrm{span}}(N_{x})\leq\theta_{\mathrm{max}}$. It can be verified that $f(x)\triangleq\operatorname{sin}\left(\operatorname{arctan}\left(x\right)-\theta_{\mathrm{max}}\right)$ is a concave increasing function with respect to $x$, and that $\lim_{x\to\infty}f^{{}^{\prime}}(x)=0$. This indicates that when $N_{x}>\bar{N}_{x}$, i.e., $\triangle_{\mathrm{span}}(N_{x})>\theta_{\mathrm{max}}$, the growth rate of the maximum SNR gradually decreases as the number of RAs further increases, eventually approaching zero.
For the infinitely large-scale ULA such that $N_{x}\to\infty$, since $\operatorname{arctan}\left(\frac{N_{x}\delta}{2}\right)\to\frac{\pi}{2}$ as $\frac{N_{x}\delta}{2}\to\infty$, the resultant SNR in (26) reduces to
@@
$\lim_{N_{x}\to\infty}\gamma_{\mathrm{fixed}}=\frac{2\xi\delta}{\pi}\bar{P},$ (28)
@@
$\frac{\lim_{N_{x}\to\infty}\gamma}{\lim_{N_{x}\to\infty}\gamma_{\mathrm{fixed}}}=\theta_{\mathrm{max}}+\operatorname{cos}\theta_{\mathrm{max}}\geq 1,$ (29)
where the inequality holds since $f(x)=x+\operatorname{cos}x$ is a monotonically increasing function with respect to $x$ and $f(0)=1$. Therefore, by exploiting the additional spatial DoFs in terms of boresight rotation to improve the array gain, the proposed RA system with the optimal pointing vectors in (20) will outperform the fixed-antenna system, and the performance gap increases with the allowable range of RA’s boresight rotation, as constrained by (4).
@@
${\tilde{\gamma}}=\begin{cases}{\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{2})-\gamma(\triangle_{1})\right],}&{\text{Case 1:}\;\phi\in\left[-\frac{\pi}{2},-\frac{N_{x}\delta}{2}\right)}\\ {\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{1})+\gamma(\triangle_{2})\right],}&{\text{Case 2:}\;\phi\in\left[-\frac{N_{x}\delta}{2},\frac{N_{x}\delta}{2}\right]}\\ {\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{1})-\gamma(\triangle_{2})\right],}&{\text{Case 3:}\;\phi\in\left(\frac{N_{x}\delta}{2},\frac{\pi}{2}\right].}\end{cases}$ (30)
In addition, since all geometric relationships shown in Fig. 4 will reduce to the symmetrical case in Fig. 3 when $N_{x}\to\infty$ regardless of $\phi$, the asymptotic SNR turns out to be the same as (27).
@@
$\left.\quad\quad\quad\int_{\operatorname{arctan}\left(\frac{D}{r}\right)-\theta_{\mathrm{max}}}^{\operatorname{arctan}\left(\frac{R}{r}\right)-\theta_{\mathrm{max}}}\operatorname{cos}^{2p}\epsilon\operatorname{tan}(\epsilon+\theta_{\mathrm{max}})\mathrm{d}\epsilon\right],$ (33)
@@
In (III-B), the first integral corresponds to the inner area with $\epsilon_{n}=0$, while the second integral corresponds to the outer area with $\epsilon_{n}>0$. The integral in (III-B) is challenging to handle since the directivity factor $p$ exists as a power exponent of the cosine function. Similar to the ULA case, we focus on the cosine gain pattern (i.e., $p=\frac{1}{2}$) for convenience in the following discussion.
Lemma 3: For $p=\frac{1}{2}$, the function $\mathcal{G}(R,p,\theta_{\mathrm{max}})$ can be expressed in closed form as (34), shown at the top of the next page, where $\theta_{R}\triangleq\operatorname{arctan}\left(\frac{R}{r}\right)$.
${\mathcal{G}\left(R,\frac{1}{2},\theta_{\mathrm{max}}\right)}=\begin{cases}{\bar{P}\xi\operatorname{ln}\left(1+\operatorname{tan}^{2}\theta_{R}\right),}&{R\leq r\operatorname{tan}\theta_{\mathrm{max}}}\\ {2\bar{P}\xi\left[1-\operatorname{ln}(\operatorname{cos}\theta_{\mathrm{max}})-\operatorname{cos}(\theta_{R}-\theta_{\mathrm{max}})+\operatorname{sin}\theta_{\mathrm{max}}\operatorname{ln}\left(\frac{(1+\operatorname{sin}\theta_{R})(1-\operatorname{sin}\theta_{\mathrm{max}})}{\operatorname{cos}\theta_{R}\operatorname{cos}\theta_{\mathrm{max}}}\right)\right],}&{R>r\operatorname{tan}\theta_{\mathrm{max}}}.\end{cases}$ (34)

### v5

@@ II-B Channel Model
where $\mathbf{F}\triangleq[\vec{\mathbf{f}}_{1},\vec{\mathbf{f}}_{2},\dots,\vec{\mathbf{f}}_{N}]\in\mathbb{R}^{3\times N}$ is the pointing matrix of all RAs, $\mathbf{h}_{k}^{\mathrm{LoS}}(\mathbf{F})\triangleq[h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{1}),h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{2}),\dots,h_{k}^{\mathrm{LoS}}(\vec{\mathbf{f}}_{N})]^{T}$ and $\mathbf{h}_{k}^{\mathrm{NLoS}}(\mathbf{F})\triangleq[h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{1}),h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{2}),\dots,h_{k}^{\mathrm{NLoS}}(\vec{\mathbf{f}}_{N})]^{T}$ are the LoS and NLoS channel components from user $k$ to the BS, respectively.
@@ II-C Min-SINR Maximization Problem
In this paper, we aim to maximize the minimum SINR among all the users by jointly optimizing the receive beamforming matrix $\mathbf{V}\triangleq[\mathbf{v}_{1},\mathbf{v}_{2},\dots,\mathbf{v}_{K}]$ and RA pointing matrix $\mathbf{F}$, subject to the rotational constraint given in (4). Thus, the min-SINR maximization problem is formulated as
$\text{(P1):}\quad\max_{\mathbf{V},\mathbf{F}}\quad$ $\min_{k}\gamma_{k}$ (14a)
@@
where constraint (14b) is equivalent to (4) for ensuring that the rotation of boresight directions does not exceed the given range, and constraint (14c) ensures that $\vec{\mathbf{f}}_{n}$ is a unit vector.
The antenna directional gain pattern in (5) can be characterized and measured after antenna fabrication and thus treated as a given scaling factor over direction. As such, the BS only needs to acquire the key channel parameters such as path loss, angle of arrival/departure (AoA/AoD), and delay for optimizing RA boresight directions. Since adjusting the antenna orientation/boresight rather than its position does not alter the propagation geometry, existing geometry-based MIMO channel estimation methods for AoA/AoD and delay estimation, such as multiple signal classification (MUSIC) and compressed sensing (CS) approaches, remain applicable under antenna rotations. Then, least-squares (LS) or minimum mean-square error (MMSE)-based estimators can be applied to further estimate the path gain. By combining these estimated parameters with the pre-measured antenna directional gain pattern, the effective channel associated with each candidate boresight direction can be efficiently constructed, thereby enabling efficient RA boresight optimization without the need to estimate the channels with different boresight directions. In addition, an efficient channel estimation scheme tailored for RA systems was proposed in [38] to further improve estimation accuracy.
Denoting $\hat{\mathbf{h}}_{k}(\mathbf{F})$ as the estimated channel state information (CSI) of the link from user $k$ to the BS, the received signal is expressed as
@@
where $\check{\mathbf{h}}_{k}\triangleq\mathbf{h}_{k}(\mathbf{F})-\hat{\mathbf{h}}_{k}(\mathbf{F})$ stands for the CSI estimation error that is uncorrelated with the estimated channel $\hat{\mathbf{h}}_{k}(\mathbf{F})$. In particular, when the number of scattering clusters is sufficiently large, both the channel and noise can be well approximated as jointly Gaussian processes. If the CSI estimation error follows the zero-mean complex Gaussian distribution, i.e., $\sum_{k=1}^{K}{\check{\mathbf{h}}_{k}\sqrt{P_{k}}s_{k}}\sim\mathcal{N}_{c}\left(\mathbf{0},e^{2}\mathbf{I}_{N}\right)$ with $e$ accounting for the level of channel estimation error [33], the resultant error term can be incorporated into the AWGN term in (15) for simplicity. Accordingly, the receive SINR for user $k$ is given by
@@ III Single-User Case with Free-Space Propagation
$\text{(P2):}\quad\max_{\mathbf{v},\mathbf{F}}\quad$ $\bar{P}|\mathbf{v}^{H}\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})|^{2}$ (17a)
@@ III-A Optimal Closed-Form Solution
For any given RA pointing matrix $\mathbf{F}$ in the single-user case, it is known that the MRC beamformer is the optimal receive beamforming solution to problem (P2) [32], i.e., $\mathbf{v}_{\mathrm{MRC}}=\frac{\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})}{\|\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})\|}$. Thus, substituting $\mathbf{v}_{\mathrm{MRC}}$ into (17a) yields the following SNR expression,
$\gamma=\bar{P}\|\mathbf{h}^{\mathrm{LoS}}(\mathbf{F})\|^{2}=\frac{\bar{P}A}{4\pi}\sum_{n=1}^{N}{\frac{G_{0}\operatorname{cos}^{2p}(\epsilon_{n})}{r_{n}^{2}}},$ (18)
where the superscript “B-U” of $r_{n}^{\mathrm{B-U}}$ is omitted in this section for simplicity. Exploiting the structure in (18), problem (P2) can be decomposed into $N$ subproblems, each of which independently optimizes the pointing vector of one RA. For RA $n$, the corresponding subproblem is given by
$\text{(P3):}\quad\max_{\vec{\mathbf{f}}_{n}}\quad$ $\operatorname{cos}(\epsilon_{n})=\vec{\mathbf{f}}_{n}^{T}\vec{\mathbf{u}}_{n}$ (19a)
@@
$\theta_{\mathrm{a},n}^{\star}$ $=\operatorname{arctan2}\left(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{2},\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{1}\right),$ (21b)
@@ III-B Asymptotic Performance Analysis
In this subsection, we focus on the performance analysis for the single-user system, and derive a closed-form expression for the SNR in the ULA case, lower/upper bounds for the SNR in the UPA case, and the asymptotic gains in both cases as the antenna number $N$ goes to infinity. For ease of exposition, we assume that the user is located along the $z$-axis (i.e., $\psi=\frac{\pi}{2}$ and $\phi=0$) and its position is denoted by $\mathbf{u}=[0,0,r]^{T}$. In this case, based on the optimal pointing vector obtained in (20), the entire array region can be divided into inner and outer areas by comparing $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})$ and $\theta_{\mathrm{max}}$, as shown in Fig. 3. Specifically, if RA $n$ is located in the inner area, we have $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})\leq\theta_{\mathrm{max}}$ and $\varpi_{n}=0$, i.e., RA $n$ can adjust its boresight to perfectly align with the user direction to obtain the maximum antenna directional gain. Conversely, for RA $n$ located in the outer area, we have $\operatorname{arccos}(\vec{\mathbf{u}}_{n}^{T}\mathbf{e}_{3})>\theta_{\mathrm{max}}$ and $\varpi_{n}=\operatorname{arctan}\left(\sqrt{(n_{x}^{2}+n_{y}^{2})\delta^{2}}\right)-\theta_{\mathrm{max}}$, i.e., RA $n$ can only serve the user with $\theta_{\mathrm{z},n}=\theta_{\mathrm{max}}$ and the RA boresight is offset from the user direction by an angle $\varpi_{n}>0$ due to the zenith angle constraint in (4). Based on the above discussion, the projection between the user direction vector and the optimal pointing vector of RA $n$ can be expressed as
$\hskip-8.5359pt\operatorname{cos}\varpi_{n}=\operatorname{cos}\left(\left[\operatorname{arctan}\left(\sqrt{(n_{x}^{2}+n_{y}^{2})\delta^{2}}\right)-\theta_{\mathrm{max}}\right]_{+}\right),$ (22)
By substituting (22) into (18), we can obtain the resultant SNR as (23), shown at the top of next page. The SNR in (23) involves a double summation, which may make it difficult to gain useful insights. By approximating the double summation in (23) as its corresponding double integral by leveraging $\delta\ll 1$ as in [15, 48, 8], the SNR can be rewritten in an integral form as (24), also shown at the top of next page.
$\gamma=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi}\sum_{n_{x}=-\frac{N_{x}-1}{2}}^{\frac{N_{x}-1}{2}}{\sum_{n_{y}=-\frac{N_{y}-1}{2}}^{\frac{N_{y}-1}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(\sqrt{(n_{x}^{2}+n_{y}^{2})\delta^{2}}\right)-\theta_{\mathrm{max}}\right]_{+}\right)}{1+(n_{x}^{2}+n_{y}^{2})\delta^{2}}}}.$ (23)
$\gamma\simeq\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi\Delta^{2}}\int_{-\frac{N_{y}\Delta}{2}}^{\frac{N_{y}\Delta}{2}}{\int_{-\frac{N_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(\sqrt{\frac{1}{r^{2}}(x^{2}+y^{2})}\right)-\theta_{\mathrm{max}}\right]_{+}\right)}{1+\frac{1}{r^{2}}(x^{2}+y^{2})}\mathrm{d}x\mathrm{d}y}}.$ (24)
@@
1) ULA-Based RA System: To gain some insights, we first focus on the ULA setting. With $N_{y}=1$ and $N_{x}=N$, the resultant SNR in (24) reduces to
$\hskip-11.38092pt\gamma=\frac{\bar{P}G_{0}\xi\delta^{2}}{4\pi\Delta}\int_{-\frac{N_{x}\Delta}{2}}^{\frac{N_{x}\Delta}{2}}{\frac{\operatorname{cos}^{2p}\left(\left[\operatorname{arctan}\left(|\frac{x}{r}|\right)-\theta_{\mathrm{max}}\right]_{+}\right)}{1+\frac{x^{2}}{r^{2}}}\mathrm{d}x}.\hskip-2.84544pt$ (25)
@@
$\hskip-5.69046pt{\gamma}\hskip-2.84544pt=\hskip-2.84544pt\begin{cases}{\frac{2\bar{P}\xi\delta}{\pi}\triangle_{\mathrm{span}}(N_{x}),}&\hskip-7.11317pt{N_{x}\leq\bar{N}_{x}}\\ {\frac{2\bar{P}\xi\delta}{\pi}\left[\theta_{\mathrm{max}}+\operatorname{sin}\left(\triangle_{\mathrm{span}}(N_{x})-\theta_{\mathrm{max}}\right)\right],}&\hskip-7.11317pt{N_{x}>\bar{N}_{x},}\end{cases}\hskip-5.69046pt$ (26)
@@
Theorem 1 shows that with the applied MRC receive beamforming and the optimized RA pointing vectors, the resultant maximum SNR of the proposed ULA-based RA system scales with the antenna number $N_{x}$ according to the span angle $\triangle_{\mathrm{span}}(N_{x})$. Furthermore, given the antenna size $A$, the antenna separation $\Delta$, and the propagation distance $r$, the maximum SNR of the ULA-based RA system depends on the allowable range for boresight rotation and the ULA size.
Remark 1: By applying the linear approximation for the arctangent function, i.e., $\operatorname{arctan}(x)\approx\frac{\pi}{4}x,\;-1\leq x\leq 1$ [23], the SNR in the first case of (26) can be approximated by $\gamma\approx\frac{1}{4}N_{x}\bar{P}\xi\delta^{2}$ since we have $0\leq\frac{N_{x}\delta}{2}\leq 1$ when $N_{x}\leq\bar{N}_{x}$. Thus, the resultant SNR increases linearly with the number of RAs when $N_{x}\leq\bar{N}_{x}$, i.e., $\triangle_{\mathrm{span}}(N_{x})\leq\theta_{\mathrm{max}}$. It can be verified that $f(x)\triangleq\operatorname{sin}\left(\operatorname{arctan}\left(x\right)-\theta_{\mathrm{max}}\right)$ is a concave increasing function, and $\lim_{x\to\infty}f^{{}^{\prime}}(x)=0$. This indicates that when $N_{x}>\bar{N}_{x}$, i.e., $\triangle_{\mathrm{span}}(N_{x})>\theta_{\mathrm{max}}$, the growth rate of the maximum SNR gradually decreases as the number of RAs further increases, eventually approaching zero.
For the infinitely large-scale ULA such that $N_{x}\to\infty$, since $\operatorname{arctan}\left(\frac{N_{x}\delta}{2}\right)\to\frac{\pi}{2}$ as $\frac{N_{x}\delta}{2}\to\infty$, the SNR in (26) reduces to
@@
$\lim_{N_{x}\to\infty}\gamma_{\mathrm{FA}}=\frac{2\xi\delta}{\pi}\bar{P},$ (28)
@@
$\frac{\lim_{N_{x}\to\infty}\gamma}{\lim_{N_{x}\to\infty}\gamma_{\mathrm{FA}}}=\theta_{\mathrm{max}}+\operatorname{cos}\theta_{\mathrm{max}}\geq 1,$ (29)
where the inequality holds since $f(x)=x+\operatorname{cos}x$ is a monotonically increasing function and $f(0)=1$. By exploiting the additional spatial DoFs in terms of boresight rotation to improve the array gain, the proposed RA system with the optimal pointing vectors in (20) will outperform the fixed-antenna system, and the performance gap increases with the allowable range of boresight rotation, as constrained by (4).
@@
$\hskip-4.26773pt{\tilde{\gamma}}\hskip-2.84544pt=\hskip-2.84544pt\begin{cases}{\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{2})-\gamma(\triangle_{1})\right],}&\hskip-7.11317pt{\text{Case 1:}\;\phi\in\hskip-2.84544pt\left[-\frac{\pi}{2},-\frac{N_{x}\delta}{2}\right)}\\ {\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{1})+\gamma(\triangle_{2})\right],}&\hskip-7.11317pt{\text{Case 2:}\;\phi\in\hskip-2.84544pt\left[-\frac{N_{x}\delta}{2},\frac{N_{x}\delta}{2}\right]}\\ {\frac{1}{2\operatorname{cos}\phi}\left[\gamma(\triangle_{1})-\gamma(\triangle_{2})\right],}&\hskip-7.11317pt{\text{Case 3:}\;\phi\in\hskip-2.84544pt\left(\frac{N_{x}\delta}{2},\frac{\pi}{2}\right].}\end{cases}\hskip-5.69046pt$ (30)
Since all geometric relationships shown in Fig. 4 will reduce to the symmetrical case in Fig. 3 when $N_{x}\to\infty$ regardless of $\phi$, the asymptotic SNR turns out to be the same as (27).
@@
$\left.\quad\quad\int_{\operatorname{arctan}\left(\frac{D}{r}\right)-\theta_{\mathrm{max}}}^{\operatorname{arctan}\left(\frac{R}{r}\right)-\theta_{\mathrm{max}}}\operatorname{cos}^{2p}\varpi\operatorname{tan}(\varpi+\theta_{\mathrm{max}})\mathrm{d}\varpi\right],$ (33)
@@
In (III-B), the first integral corresponds to the inner area with $\varpi_{n}=0$, while the second integral corresponds to the outer area with $\varpi_{n}>0$. The integral in (III-B) is challenging to handle since the directivity factor $p$ exists as a power exponent of the cosine function. Similar to the ULA case, we focus on the cosine gain pattern (i.e., $p=\frac{1}{2}$) for convenience in the following discussion.
Lemma 3: For $p=\frac{1}{2}$, the function $\mathcal{G}(R,p,\theta_{\mathrm{max}})$ can be expressed in closed form as (34), shown at the top of next page, where $\theta_{R}\triangleq\operatorname{arctan}\left(\frac{R}{r}\right)$.
$\hskip-2.84544pt{\mathcal{G}\left(R,\frac{1}{2},\theta_{\mathrm{max}}\right)}=\begin{cases}{\bar{P}\xi\operatorname{ln}\left(1+\operatorname{tan}^{2}\theta_{R}\right),}&{R\leq r\operatorname{tan}\theta_{\mathrm{max}}}\\ {2\bar{P}\xi\left[1-\operatorname{ln}(\operatorname{cos}\theta_{\mathrm{max}})-\operatorname{cos}(\theta_{R}-\theta_{\mathrm{max}})+\operatorname{sin}\theta_{\mathrm{max}}\operatorname{ln}\left(\frac{(1+\operatorname{sin}\theta_{R})(1-\operatorname{sin}\theta_{\mathrm{max}})}{\operatorname{cos}\theta_{R}\operatorname{cos}\theta_{\mathrm{max}}}\right)\right],\hskip-5.69046pt}&{R>r\operatorname{tan}\theta_{\mathrm{max}}}.\end{cases}$ (34)